# StreetCrafter: Street View Synthesis with Controllable Video Diffusion Models

### v1

@@ Abstract
footnotetext: Equal contribution.
@@ 2 Related Work
To better support downstream applications like reconstruction, some works have proposed camera-controlled video generation methods. [70, 21, 65] propose training-free methods that control the denoising process of an existing diffusion model to achieve controllable video generation. While these methods do not require training or fine-tuning the diffusion model, their performance is limited due to ambiguity during generation. Some recent works finetune the video diffusion model with additional camera parameters as additional input. [55, 26, 34] generate novel views for object-centric scene. Some methods achieve camera control by inject camera parameters into the video diffusion model [66, 60, 61, 1, 54] based on the architecture of U-Net or transformer [40]. Another line of works use geometry foundation models to build explicit representation such as point cloud as guidance to the video diffusion model [29, 71, 36]. However, they mainly focus on static scenes while our method leverages the more accurate LiDAR prior to provide guidance for complex driving scenes.
@@
There have been works that try to utilize the LiDAR input [4, 39, 49, 33, 48] to enhance the model’s capability of capturing scene geometry and generalizing to novel viewpoints. However, they mainly deal with static scenes only using supervision from the input trajectory while our model combines the LiDAR with generative prior to generate guidance on novel trajectories for dynamic urban scenes.
Reconstruction with Diffusion Prior  [28, 41] use SDS for lifting 2D generation to 3D, achieving text-based 3D generation. [30, 27, 32] generate multi-view predictions based on a single-view image input using diffusion model, and then use multi-view reconstruction methods to obtain the reconstructed 3D model. [62] enhances sparse-view reconstruction using diffusion prior, and [11] leverages multi-view diffusion to improve multi-view consistency. For scenes with a certain scale, video generation models can more efficiently generate a larger number of novel views with great multi-view consistency compared to image generation models. [29, 71] use DUSt3R [57] to reconstruct the point cloud from sparse viewpoints, leveraging the geometry structure as a condition to guide the video generation model in producing novel-view images, enabling the reconstruction of a static scene with minimal viewpoint input. [6, 31] utilize the degraded rendering results as condition to perform sparse view synthesis with video diffusion models. Recent works [74, 13, 9, 77] enhance street scene reconstruction quality by incorporating diffusion prior to generate additional views, similar to our methods. However, their diffusion models lack precise camera control, which limits the accuracy and visual quality of their reconstruction.
@@ 5.1 Experiments Setup
We evaluate the performance of StreetCrafter and the crafted dynamic 3DGS representation which we referr to as Ours-V and Ours-G respectively on the task of novel view synthesis. For Ours-V, we crop and resize the input image to 576 $\times$ 1024 to match the output resolution of video diffusion model during evaluation. Datasets.  We conduct experiments on Waymo Open Dataset [47] and PandaSet [64], using their 10Hz front camera and synchronized LiDAR. We select 15 sequences of approximately 100 frames from the validation set of Waymo and 5 sequences of 80 frames from PandaSet to test the novel view synthesis results. We uniformly sample half of the images in each sequence as the testing frames and use the remaining for training. Input image resolution is set to $1066\times 1600$ and $900\times 1600$ for Waymo and PandaSet, respectively. The training set of Waymo and the remaining PandaSet sequences are used to train StreetCrafter, resulting in a total of approximately 35,000 training samples. More details can be found in the supplementary.
@@ 6 Conclusion
Appendix A More Implementation Details
A.1 StreetCrafter Training Details
@@
For the adaptation from the pretrained model of Vista [12], we ignore the action control layer injected via cross-attention and mark the first element of the frame-wise mask to 1 and the rest to 0. We incorporate the LoRA [22] adapters introduced during the learning of action controllability as it contributes to the enhancement of visual quality [12]. More details can be found in the original paper.
@@
A.2 StreetCrafter Distillation Details
Loss function.
@@ A.2 StreetCrafter Distillation Details
We jointly optimize the gaussian parameters of background and foreground moving objects, texel of the high-resolution sky cubemap and noisy object tracklets following Street Gaussians [67]. The extra loss $\mathcal{L}_{\text{g}}$ for input view camera is defined as:
$\mathcal{L}_{\text{g}}=\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{sky}}\mathcal{L}_{\text{sky}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}.$ (10)
For $\mathcal{L}_{\text{depth}}$, we apply the $\mathcal{L}_{1}$ loss between rendered depth $\mathbf{D}$ and the LiDAR measurement’s depth $\mathbf{D}^{\text{lidar}}$ where we optimize 95% of the pixels with smallest depth error:
$\mathcal{L}_{\text{depth}}=\sum||\mathbf{D}-\mathbf{D}^{\text{lidar}}||_{1}.$ (11)
For $\mathcal{L}_{\text{sky}}$, we apply the binary cross entropy loss between rendered opacity $\mathbf{O}_{g}$ and predicted sky mask $\mathbf{M}_{\text{sky}}$ generated by Grounded SAM [43]:
$\mathcal{L}_{\text{sky}}=-\sum((1-\mathbf{M}_{\text{sky}})\text{log}\mathbf{O}_{g}+\mathbf{M}_{\text{sky}}\text{log}(1-\mathbf{O}_{g})).$ (12)
For $\mathcal{L}_{\text{reg}}$, the regularization term in our loss function is defined as an entropy loss on the accumulated alpha values of decomposed foreground objects $\mathbf{O}_{\text{obj}}$:
$\mathcal{L}_{\text{reg}}=-\sum(\mathbf{O}_{\text{obj}}\text{log}\mathbf{O}_{\text{obj}}+(1-\mathbf{O}_{\text{obj}})\text{log}(1-\mathbf{O}_{\text{obj}})).$ (13)
@@
The coefficients $\lambda_{\text{depth}}$, $\lambda_{\text{sky}}$ and $\lambda_{\text{reg}}$ in Equation 10 are set to 0.01, 0.05 and 0.1, respectively. For the loss function of novel view camera, we crop the upper part of the rendering image and resize to $576\times 1024$ to compute the LPIPS [76] loss with novel view image generated by StreetCrafter .
Point cloud initialization.
@@
We initialize the background gaussian model as the combination of LiDAR and SfM point cloud following. The object gaussian model is initialized with aggregated LiDAR points obtained from object tracklets or random sampling. The colors of LiDAR points are assigned by projecting them onto the image plane.
Optimization.
@@
We adopt the densification strategy introduced in [73] to prevent suboptimal solutions by accumulating the norms of view-space position gradients. The densification threshold is set to 0.0006. We disable the pruning of big gaussians in world space since this hinders the gaussian model to represent distant regions and the LiDAR points have provided a good initialization to prevent the model from falling into local optima. We finally introduce the 2D Mip filter to enable anti-aliased rendering inspired by [72].
@@
We sample StreetCrafter every 5000 iterations from iteration 7000 to 22000 and linearly reduce the noise scale $s$ from 0.7 to 0.3. We use the annotated object tracklets provided by the datasets, with the learning rates for the translation vector and rotation matrix initialized at $5e^{-4}$ and $1e^{-5}$, respectively, decaying exponentially to $1e^{-5}$ and $5e^{-6}$. For the remaining parameters, we use the default values from the official implementation of Street Gaussians [67].
A.3 Evaluation Details
Interpolation of StreetCrafter.
@@ A.3 Evaluation Details
For the interpolation setting of Ours-V in the main paper, we incorporate training images along the input trajectory in addition to the reference image and LiDAR conditions. During each denoising step, we replace the prediction of $\mathcal{F}_{\theta}$ at training frame with the clean latent of training images. This could lead to improvement in the interpolation quality, with PSNR increasing from 23.66 to 27.19 and LPIPS decreasing from 0.098 to 0.087 on Waymo Open Dataset [47].
Baselines.
@@
We use the same object tracklets as our method for all the baselines requiring 3D bounding box as input [67, 52, 69]. We use the same rendering kernel and optimization strategy as our method for all the baselines using 3DGS as the scene representation [24, 67].
Appendix B Additional Experiments
B.1 More Editing Results
@@
B.2 More Comparisons
We provide more qualitative comparisons on Waymo [47] dataset under the setting of lane change in Figure 10.
B.3 Interpolation Results
Figures 11, 12 display the view interpolation results on Waymo Open [47] and PandaSet [64] datasets. Our method achieves comparable rendering quality to the baselines while delivering significantly better results for view extrapolation.
B.4 More Ablations

### v2

@@ Abstract
footnotetext: Equal contribution. Corresponding author.
@@ 2 Related Work
To better support downstream applications such as reconstruction, some works have proposed camera-controlled video generation methods. [71, 21, 66] propose training-free methods that control the denoising process of an existing diffusion model to achieve controllable video generation. While these methods do not require training or fine-tuning the diffusion model, their performance is limited due to ambiguity during generation. Some recent works finetune the video diffusion model with additional camera parameters as additional input. [56, 27, 36] generate novel views for object-centric scene. Some methods achieve camera control by injecting camera parameters into the video diffusion model [67, 61, 62, 1, 55] based on the architecture of U-Net or transformer [42]. Another line of works use geometry foundation models [58, 30] to build explicit representation such as point cloud as guidance to the video diffusion model [31, 72, 38]. However, they mainly handle static scenes, while our method leverages the more accurate LiDAR prior to control complex driving scenes.
@@
There have been works that try to utilize the LiDAR input [4, 41, 50, 35, 49] to enhance the model’s capability of capturing scene geometry and generalizing to novel viewpoints. However, they mainly deal with static scenes only using supervision from the input trajectory, while our model combines the LiDAR with generative prior to generate guidance on novel trajectories for dynamic urban scenes.
Reconstruction with Diffusion Prior  [29, 43] use SDS for lifting 2D generation to 3D, achieving text-based 3D generation. [32, 28, 34] generate multi-view predictions based on a single-view image input using diffusion model, and then use multi-view reconstruction methods to obtain the reconstructed 3D model. [63, 11] further extend the diffusion prior to scene level. Some works have utilized video diffusion models to generate novel views with improved multi-view consistency compared to image diffusion models. [31, 72] employ DUSt3R [58] to build coarse geometry structure from sparse viewpoints, which serves as guidance for the video diffusion model, allowing it to produce additional inputs for reconstructing the static scene. [6, 33] use the degraded rendering results as condition to perform sparse view synthesis with video diffusion models. Recent works [75, 13, 9, 78, 23] enhance street scene reconstruction quality by incorporating diffusion prior to generate additional views, similar to our methods. However, their diffusion models lack precise camera control, which limits the accuracy and visual quality of their reconstruction.
@@ 5.1 Experiments Setup
We evaluate the performance of StreetCrafter and the crafted dynamic 3DGS representation which we refer to as Ours-V and Ours-G respectively on the task of novel view synthesis. For Ours-V, we crop and resize the input image to 576 $\times$ 1024 to match the output resolution of video diffusion model during evaluation. Datasets.  We conduct experiments on Waymo Open Dataset [48] and PandaSet [65], using their 10Hz front camera and synchronized LiDAR. We select 15 sequences of approximately 100 frames from the validation set of Waymo and 5 sequences of 80 frames from PandaSet to test the novel view synthesis results. We uniformly sample half of the images in each sequence as the testing frames and use the remaining for training. Input image resolution is set to $1066\times 1600$ and $900\times 1600$ for Waymo and PandaSet, respectively. The training set of Waymo and the remaining PandaSet sequences are used to train StreetCrafter, resulting in a total of approximately 35,000 training samples. More details can be found in the supplementary.
@@ 6 Conclusion
Acknowledgement.  The authors would like to acknowledge the support from NSFC (No. 62402427, U24B20154), Li Auto, Information Technology Center and State Key Lab of CAD&CG, and Zhejiang University Education Foundation Qizhen Scholar Foundation.
@@
7 More Implementation Details
7.1 StreetCrafter Training Details
@@
For adaptation from the pretrained model of Vista [12], we ignore the action control layer injected via cross-attention and mark the first element of the frame-wise mask to 1 and the rest to 0. We incorporate the LoRA [22] adapters introduced during the learning of action controllability as it contributes to the enhancement of visual quality [12]. More details can be found in the original paper.
@@
7.2 StreetCrafter Distillation Details
@@
Loss function.  We jointly optimize the gaussian parameters of background and foreground moving objects, texel of the high-resolution sky cubemap and noisy object tracklets following Street Gaussians [68]. The extra loss $\mathcal{L}_{\text{g}}$ for input view camera is defined as:
$\mathcal{L}_{\text{g}}=\lambda_{\text{depth}}\mathcal{L}_{\text{depth}}+\lambda_{\text{sky}}\mathcal{L}_{\text{sky}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (10)
@@
where $\mathcal{L}_{\text{depth}}$, $\mathcal{L}_{\text{sky}}$ and $\mathcal{L}_{\text{reg}}$ share the same format as Street Gaussians. Please refer to the original paper for more details. The coefficients $\lambda_{\text{depth}}$, $\lambda_{\text{sky}}$ and $\lambda_{\text{reg}}$ in Equation 10 are set to 0.01, 0.05 and 0.1, respectively. For the loss function of novel view camera, we crop the upper part of the rendering image and resize to $576\times 1024$ to compute the LPIPS [77] loss with novel view image generated by StreetCrafter .
@@
Point cloud initialization.  We initialize the background gaussian model as the combination of LiDAR and SfM point cloud. The object gaussian model is initialized with aggregated LiDAR points obtained from object tracklets or random sampling. The colors of LiDAR points are assigned by projecting them onto the nearest image plane.
@@
Optimization.  We adopt the densification strategy introduced in [74] to prevent suboptimal solutions by accumulating the norms of view-space position gradients. The densification threshold is set to 0.0006. We disable the pruning of big gaussians in world space since this hinders the gaussian model to represent distant regions and the LiDAR points have provided a good initialization to prevent the model from falling into local optima. We finally introduce the 2D Mip filter to enable anti-aliased rendering inspired by [73].
@@
We sample StreetCrafter every 5000 iterations from iteration 7000 to 22000 and linearly reduce the noise scale $s$ from 0.7 to 0.3. We use the annotated object tracklets provided by the datasets, with the learning rates for the translation vector and rotation matrix initialized at $5e^{-4}$ and $1e^{-5}$, respectively, decaying exponentially to $1e^{-5}$ and $5e^{-6}$. For the remaining parameters, we use the default values from the official implementation of Street Gaussians.
7.3 Evaluation Details
@@
Interpolation.  For the interpolation setting of Ours-V in the main paper, we incorporate training images along the input trajectory in addition to the reference image and LiDAR conditions. During each denoising step, we replace the prediction of $\mathcal{F}_{\theta}$ at training frame with the clean latent of training images. This could lead to improvement in the interpolation quality, with PSNR increasing from 23.66 to 27.19 and LPIPS decreasing from 0.098 to 0.087 on Waymo Open Dataset [48].
@@
Baselines.  We use the same object tracklets as our method for all the baselines requiring 3D bounding box as input [68, 53, 70]. We use the same rendering kernel and optimization strategy as our method for all the baselines using 3DGS as the scene representation [25, 68].
Metric.  For Fréchet inception distance (FID) metric, we mark the input video as real and the rendered sequence as unreal for each scene. For Ours-V, we upper-crop the input video to match the resolution of the generated video.
8 Additional Experiments
8.1 More Comparisons
Comparisons with baselines.  We provide more qualitative comparisons on Waymo [48] dataset under the setting of lane change in Figure 13. Figures 14, 15 display the view interpolation results on Waymo Open [48] and PandaSet [65] datasets. Our method achieves comparable rendering quality to the baselines while achieving significantly better results for view extrapolation.
Comparisons with concurrent works.  We compare Ours-V with ViewCrafter [72] and Ours-G with DriveDreamer4D [78], both of which are concurrent of our works. For ViewCrafter, we make several modifications to improve its performance. First, we build the global point cloud from the whole input sequence instead of selecting the first frame as in the original setting. Second, we fix the camera parameters during the global alignment process of DUSt3R [58] by using camera calibration results from the dataset. This can also help define camera poses within the dataset’s coordinate system when performing view extrapolation. For DriveDreamer4D, we compare our method on segment-103593 of Waymo following their setting with PVG [5] as the base model.
As shown in Table 5 and Figures 8, 9, our method achieves better view synthesis results under both input and novel trajectory compared with ViewCrafter [72]. ViewCrafter builds point cloud without considering that the geometry of dynamic scene changes overtime, thus it fails to accurately model dynamic regions as shown in Figure 9. The generated results also degrade significantly as the camera deviates from the input trajectory since the predicted point cloud is defined in the camera coordinate system. In contrast, the LiDAR point cloud defined in the world coordinate system provides our model with stronger generalization ability on new trajectories even if no ground truth data is available during training.
As shown in Figure 10, our method achieves better results under novel trajectory compared with DriveDreamer4D [78] (FID @ 2m 89.71 vs. 91.23, FID @ 3m 96.11 vs. 123.32). Due to the sparse conditioning of DriveDreamer4D, the generated videos often lack accurate 3D perception. This leads to noticeable artifacts in the reconstructed scene, such as the vehicle in the lower left corner and the building on the right side.
Methods Input trajectory Novel trajectory
[…]
ViewCrafter [72] 21.59 0.226 97.86 135.69 137.76
Ours-V 25.90 0.143 60.49 62.43 73.49
Table 5: Quantitative comparison with ViewCrafter [72]. We use the video clips in ablation to test the results on input trajectory and the scenes in experiment to test the results on novel trajectory. Metrics are averaged over all sampled sequences.
Figure 8: Qualitative comparison with ViewCrafter [72] under novel trajectory. The camera is laterally shifted for 3 meters. Input view denotes the closest input video frames.
Figure 9: Qualitative comparison with ViewCrafter [72] under input trajectory. Our model can handle moving objects.
Figure 10: Qualitative comparison with DriveDreamer4D [78] under novel trajectory. The camera is laterally shifted for 2 meters.
8.2 More Editings
@@
8.3 More Ablations